# Commissioning of a radiofrequency quadrupole cooler-buncher for collinear laser spectroscopy

### v1

@@ III Performance Test
Figure 4: (a) Ion count rate and width of the time of flight spectrum of the 85Rb ion bunches from the RFQ cooler-buncher system as a function of the buffer gas flow rate. (b) Ion count rate of the 85Rb ion bunches as a function of the voltage amplitude of the applied RF signal at different RF frequencies.
Figure 5: The time of flight (TOF) spectra of stable 85Rb and 115In ions detected by the MagneTOF ion detector, with the RFQ cooler-buncher operation turned on and off. (a) TOF spectrum of 85Rb produced from the surface ion source. (b) TOF spectrum of 115Rb produced from the laser ablation ion source.
Table 1: A summary of the optimized operation parameters for the RFQ cooler-buncher system, tested with the Rb ions from the surface ion source. All values are directly read from device, except for DC potential gradient along the axial direction, which is obtained based on the COMSOL simulation using set voltages.
@@
Buffer Gas Flowrate (SCCM) 20 RF Frequency (kHz) 500
@@
The RFQ cooler-buncher system described above has been successfully manufactured and installed at Peking University. This system will be used to generate bunched ion beams for subsequent collinear resonance laser spectroscopy experiments. Thus, our primary focus is the TOF spectrum width of the ion bunches and its overall efficiency. In order to evaluate performance, the bunched ions are detected by a MagneTOF ion detector (ETP, 14924 MagneTOFTM Mini) located downstream in the laser spectroscopy setup, as shown in Fig.1 of Ref. CRIS-NIMB-PKU . The ion bunches generated from the cooler-buncher are delivered towards the ion detector through a 90 degree dipole magnet, as depicted in Fig. 1.
Given the large number of parameters involved in the RFQ cooler-buncher system operation, a systematic testing approach is required, including the testing of all the DC potentials, RF amplitude and frequency, buffer gas flowrate, and other factors. Prior to the experimental tests, simulation of this system was performed using COMSOL Multiphysics to identify suitable initial parameters for the testing process. Starting from these initial values, each component is tested and individually optimized, followed by iterative fine-tuning of various devices to achieve global optimization. For example, as shown in Fig. 4, the effects of the buffer gas flowrate and RF parameters on the ion bunches count rate and TOF width are systematically examined. Using stable Rb ions from the surface ion source, the maximum count rate of the ion bunches is achieved with an RF voltage of $U_{\rm{RF}}=167$ V, an RF frequency of $f=500$ kHz, and a helium flowrate of 20 SCCM, as shown in Fig. 4. Notably, throughout this optimization process, the width of TOF spectrum of ion bunches is maintained at approximately 2 $\mu$s, a value for this key parameter that is suitable for laser spectroscopy experiments.
@@
Other parameters can also influence the properties of the bunched ion beam. For example, the DC- and DC+ voltages applied to the extraction endcap significantly affect both the ion count rate and the TOF spectrum. And the potential gradient across the 12 DC electrodes will change the profile of TOF spectrum (e.g., symmetrical vs. asymmetrical). In principle, a TOF spectrum with a width as narrow as few tens of nanoseconds can be achieved, as demonstrated at MIRACLS Carina2025 , but this comes at the cost of an increased energy spread and possibly a loss in overall efficiency. However, considering the commonly used TOF width of approximately 2 $\mu$s for collinear laser spectroscopy 79Zn ; CRIS-Fr , the optimized parameters for a transmission of 60% Rb ions are listed in Tab.1. Under these conditions, the observed TOF spectra of 85Rb ions from the surface ion source and 115In ions from a laser ablation ion source are shown in Fig.5. A comparison of the TOF spectra with and without RFQ cooler-buncher operation clearly demonstrates that this system effectively compresses the continuous ion beam and the bunched ion beam with a large temporal width of $\sim$100 $\mu$s, narrowing it into a significantly smaller ion bunches with a temporal width of $\sim$ 2 $\mu$s, while maintaining an overall efficiency greater than 60%.
@@
Our ultimate goal is to integrate the RFQ cooler-buncher system into the newly-developed collinear resonance laser spectroscopy setup CRIS-NIMB-PKU , aimed at measuring the nuclear properties of unstable nuclei far from $\beta$-stability on the nuclear chart. However, as discussed in Ref. CRIS-NIMB-PKU , the earlier test of collinear resonance ionization laser spectroscopy, without the RFQ cooler-buncher system, encountered a major issue: an unacceptable overall ion beam transmission efficiency of less than 40%. This problem primarily arises from the distorted shape of the ion beam delivered into the CLS beamline, which significantly impacted the collinear overlap between the ion beam and laser beams within the meter-long interaction region, thereby reducing the resonance ionization spectroscopy efficiency PengZhangPhD . With the successful integration of the RFQ cooler-buncher between the ion source and the laser spectroscopy setup, we have now achieved an optimal overall ion beam transmission efficiency of $>$80%, from FC2 to the end section of the CLS beamline, as shown in Fig.1 of Ref. CRIS-NIMB-PKU . This demonstrates that, in addition to producing shorter ion bunches, the RFQ cooler-buncher system also improves the ion beam profile by reducing the beam’s transverse emittance.
@@ IV Conclusion and Near term experiment plan
In summary, a RFQ cooler-buncher system and its commissioning test are presented. The results demonstrate that this system successfully fulfills its intended functions of compressing the ion beam into short bunches and improving the ion beam profile. With an efficiency of RFQ cooler-buncher exceeding 60%, ion beam bunches with a temporal width of $\sim$2 µs are achieved. Furthermore, the integration of this newly-developed system into the collinear laser spectroscopy setup has largely improved the overall transmission efficiency of the ion beam due to the enhanced beam profile.
Building on these commissioning results, the RFQ cooler-buncher system will be firstly used for the offline test of a collinear resonance ionization laser spectroscopy setup CRIS-NIMB-PKU and then for online laser spectroscopy experiment of unstable nuclei at BRIF facility BRIF targeting Rb and Cs isotopes. Using a UCX target, neutron-rich 85-100Rb and 147-150Cs isotopes have already been produced at BRIF facility with a yield exceeding 100 particles per second. Given the relatively large energy spread of the ion beam at BRIF, approximately 21 eV for 38K as observed in an earlier experiment WANG2022166622 , the integration of the RFQ cooler-buncher system will be crucial not only for producing bunched ion beams but also for reducing the energy spread of the radioactive ion beam. This will undoubtedly enhance the sensitivity and spectral resolution of the planned online collinear resonance ionization laser spectroscopy experiments.

### v3

@@ III Performance Test
Figure 3: (a) Count rate of the 85Rb ion bunches as a function of the RF voltage amplitude at different RF frequencies. (b) Counts per bunch and TOF width as functions of the RFQ trapping time.
Table 1: A summary of the optimized operating parameters for the RFQ cooler-buncher system, tested with Rubidium ions from the surface ion source. All values are directly read from the device, except for the DC potential gradient along the axial direction, which is calculated based on the voltages applied to the 12 DC electrodes.
@@
Trapping time (ms) 5 RF Frequency (kHz) 500
Figure 4: The electrostatic potential along the RFQ axial direction calculated using voltages listed in Table 1. The red line indicates the potential in the trapping mode, while the blue line corresponds to the extraction mode.
Figure 5: TOF spectra of stable 85Rb and 115In ions detected by the MagneTOF ion detector, with the RFQ cooler-buncher operation off and on. (a) TOF spectra of 85Rb produced from the surface ion source. (b) TOF spectra of 115In produced from the laser ablation ion source.
Figure 6: The measured color-coded two-dimensional spectrum of TOF versus frequency for 85Rb (see Ref. Hu et al. (2025) for details). The projections onto the $x$-axis and $y$-axis correspond to the HFS spectrum (a) and TOF spectrum (b), respectively.
@@
The RFQ cooler-buncher system described above has been successfully manufactured and installed at Peking University. Since this system will be used to generate bunched ion beams for subsequent collinear resonance laser spectroscopy experiments, our primary focus is the TOF spectrum width of the ion bunches and overall transmission efficiency. In order to evaluate performance, the bunched ions are detected by a MagneTOF ion detector (ETP, 14924 MagneTOFTM Mini) located downstream in the laser spectroscopy setup, as shown in Fig. 1 of Ref. Zhang et al. (2023). The ion bunches generated from the cooler-buncher are delivered towards the ion detector through a 90 degree dipole magnet, as depicted in Fig. 1.
Given the large number of parameters involved in the RFQ cooler-buncher system operation, a systematic testing approach is required. During commissioning, each component of the RFQ system is individually tested and optimized, followed by iterative fine-tuning of various device parameters to achieve global optimization. For example, as shown in Fig. 3, the influence of the RF parameters and trapping time on the released ion counts and TOF width is examined. Using stable Rb ions from the surface ion source, the maximum count rate is achieved with an RF voltage of $U_{\rm{RF}}=167$ V and an RF frequency of $f=500$ kHz. With regard to trapping time, when ions are confined in the RFQ for less than 1 ms, the duration is insufficient for effective cooling and trapping, resulting in reduced counts per bunch and a broader TOF width. As the trapping time increases, the released counts per bunch reach a maximum and remain on a plateau within the range of 1–100 ms. Beyond 100 ms, a gradual decline in counts is observed, suggesting potential ion losses during prolonged confinement. As shown in Fig. 3(b), except for trapping time below 1 ms, the TOF width of the ion bunch remains constant at approximately 2 $\mu$s. Taking both the counts per bunch and the TOF width into consideration, a trapping time of 5 ms is adopted during commissioning. The RFQ is operated at a repetition rate of 100 Hz to enable synchronization with the 100 Hz laser pulses used in the subsequent collinear resonance ionization spectroscopy measurement Hu et al. (2025).
@@
Other parameters can also influence the properties of the bunched ion beam. For example, the DC- and DC+ voltages applied to the extraction endcap significantly affect both the ion count rate and the TOF spectrum. The potential gradient across the 12 DC electrodes will change the profile of the TOF spectrum (e.g., symmetrical or asymmetrical). The buffer gas flow rate in the range of 10–30 SCCM has only a minor effect on both the ion count rate and TOF width. In principle, a TOF spectrum with a width as narrow as few tens of nanoseconds can be achieved, as demonstrated at MIRACLS, but this comes at the cost of an increased energy spread and possibly a loss in overall efficiency. However, considering the commonly used TOF width of approximately 2 $\mu$s for collinear laser spectroscopy Yang et al. (2016); Lynch et al. (2014), the optimized parameters listed in Table 1 are obtained under conditions that yield an overall RFQ transmission efficiency of 60% for both continuous mode and bunched mode, as measured by using the FC1 and FC2. The electrostatic potential along the axial direction of the RFQ cooler-buncher, calculated using the actual voltage settings listed in Table 1, is shown in Fig. 4. Under these conditions, the observed TOF spectra of 85Rb ions from the surface ion source and 115In ions from a laser ablation ion source are shown in Fig. 5. A comparison of the TOF spectra with and without RFQ cooler-buncher operation clearly demonstrates that this system effectively compresses the continuous ion beam and the bunched ion beam with a large temporal width of $\sim$100 $\mu$s, narrowing it into a significantly narrower ion bunches with a temporal width of $\sim$2 $\mu$s, while maintaining an overall efficiency greater than 60%. In addition, the RFQ cooler-buncher demonstrated consistent and reliable performance during both this test and the laser spectroscopy experiment Hu et al. (2025), operating continuously for 8 to 10 hours per day while providing sustainable beam quality without noticeable degradation. This further confirms the long-term operational reliability of the RFQ system.
@@
The goal is to integrate the RFQ cooler-buncher system into the collinear resonance laser spectroscopy setup Zhang et al. (2023), aimed at measuring the nuclear properties of unstable nuclei far from $\beta$-stability on the nuclear chart. However, as discussed in Ref. Zhang et al. (2023), the earlier test of collinear resonance ionization laser spectroscopy, without the RFQ cooler-buncher system, encountered a major issue: an unacceptable overall ion beam transmission efficiency of less than 40%. This problem primarily arose from the distorted shape of the ion beam delivered into the CLS beamline, which significantly impacted the overlap between the ion beam and laser beams within the meter-long interaction region, thereby reducing the resonance ionization efficiency Zhang (2024). With the successful integration of the RFQ cooler-buncher between the ion source and the laser spectroscopy setup, we have now achieved an overall ion beam transmission efficiency of $>$80%, from FC2 to the end section of the CLS beamline, as shown in Fig. 1 of Ref. Zhang et al. (2023). This demonstrates that, in addition to producing narrower ion bunches, the RFQ cooler-buncher system also improves the ion beam profile by reducing its transverse emittance, which is crucial for achieving optimal laser-ion overlapping during the CLS experiment Hu et al. (2025). On this basis, we have measured the high-resolution and high efficiency HFS spectrum of stable 85Rb, as shown in Fig. 6 (see Ref. Hu et al. (2025) for details). These results confirm that the RFQ has successfully fulfilled its intended functions for laser spectroscopy experiments.
@@ IV Conclusion and Near term experiment plan
In summary, an RFQ cooler-buncher system and its commissioning test are presented. The results demonstrate that this system successfully fulfills its intended functions of compressing the ion beam into short bunches and improving the ion beam profile. With an efficiency of RFQ cooler-buncher exceeding 60%, ion beam bunches with a temporal width of $\sim$2 µs are achieved. Furthermore, the integration of this newly-developed system into the offline collinear laser spectroscopy setup has largely improved the overall transmission efficiency of the ion beam due to the enhanced beam profile with a lower transversal emittance.
Building on these commissioning results, the RFQ cooler-buncher system is firstly used for the offline test of a collinear resonance ionization laser spectroscopy setup Hu et al. (2025). Soon it will be integrated into the online laser spectroscopy experiment of unstable nuclei at BRIF facility Nan et al. (2025) targeting Rb and Cs isotopes. Using a heavy-element target, neutron-rich 85-100Rb and 147-150Cs isotopes have already been produced at BRIF facility with a yield exceeding 100 particles per second. Given the relatively large energy spread of the ion beam at BRIF, approximately 21 eV for 38K as observed in an earlier experiment Wang et al. (2022), the installation of the RFQ cooler-buncher system will be crucial not only for producing bunched ion beams but also for reducing the energy spread of the radioactive ion beam. This will undoubtedly enhance the sensitivity and spectral resolution of the planned online collinear resonance ionization laser spectroscopy experiments.